What is the avg variable cost of the second unit produces?

Quantity total cost

0 20

1 25

2 35

3 50


1
Expert's answer
2021-08-03T11:50:18-0400

Total Cost is the sum of fixed cost and variable costs.

Total Cost=Total Fixed Cost+Total Variable Cost

Total Fixed cost is that cost when quantity of the output is 0. In the table, we can see that when output is zero, Total Cost is 20.

So, "TFC=20"

From the next output level, we can find out "TVC=TC-TFC"

Average Variable Cost"=\frac{Total\space Variable\space Cost}{Output}"




For the second unit,

"TC=35"

"TVC=35-20=15"

"TVC=\frac{15}{2}=7.5"

So, the average variable cost of the second unit produces=7.5